### v7

# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

Abstract
Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years. In this paper, we introduce a more efficient gradient variant, called quadratic gradient, for privacy-preserving logistic regression training. We enhance Nesterov’s Accelerated Gradient (NAG) and Adaptive Gradient Algorithm (Adagrad) by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets. Experiments show that the enhanced algorithms have state-of-the-art performance in terms of convergence speed when compared to their raw first-order gradient methods. Also, we employ the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results in only 4 iterations. There is a promising possibility that quadratic gradient could combine first-order gradient descent/ascent algorithms with the second-order Newton-Raphson method, making it applicable to a wide range of numerical optimization problems.
Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
1 Introduction
Given a person’s healthcare data related to a certain disease, we can train a logistic regression (LR) model capable of telling whether or not this person is likely to develop this disease. However, such personal health information is highly private to individuals. The privacy concern, therefore, becomes a major obstacle for individuals to share their biomedical data. The most secure solution is to encrypt the data into ciphertexts first by Homomorphic Encryption (HE) and then securely outsource the ciphertexts to the cloud, without allowing the cloud to access the data directly. iDASH is an annual competition that aims to call for implementing interesting cryptographic schemes in a biological context. Since 2014, iDASH has included the theme of genomics and biomedical privacy. The third track of the 2017 iDASH competition and the second track of the 2018 iDASH competition were both to develop homomorphic-encryption-based solutions for building an LR model over encrypted data.
Several studies on logistic regression models are based on homomorphic encryption. Kim et al. (2018b) discussed the problem of performing LR training in an encrypted environment. They used the full batch gradient descent in the training process and the least-squares method to get the approximation of the sigmoid function. In the iDASH 2017 competition, Bonte and Vercauteren (2018), Kim et al. (2018a), Chen et al. (2018), and Crawford et al. (2018) all investigated the same problem that Kim et al. (2018b) studied. In the iDASH competition of 2018, Kim et al. (2019b) and Blatt et al. (2019) further worked on it for an efficient packing and semi-parallel algorithm. There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a). Bonte and Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method. Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al. (2018a) for efficient homomorphic computation.
Our specific contributions in this paper are as follows:
  1. 1.
    We propose a new gradient variant, quadratic gradient, which can combine the first-order gradient algorithms and the second-order Newton-Raphson method (aka Newton’s method) as one.
  2. 2.
    We develop two enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms exhibit state-of-the-art performance in terms of convergence speed.
  3. 3.
    We implement privacy-preserving logistic regression training using the enhanced NAG method, to our best knowledge, which seems to be a good choice without compromising much on computation and storage.
2 Preliminaries
We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows. For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
2.1 Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (2009). FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext (Jäschke & Armknecht, 2016) also contributes to the slowdown. Cheon et al. (2017) proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a, b; Han et al., 2019).
2.2 Database Encoding Method
Kim et al. (2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources. Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
When employing this encoding scheme, we can manipulate the data matrix $Z$ through HE operations on the ciphertext $Enc[Z]$, utilizing only three HE operations - rotation, addition, and multiplication. For instance, if we wish to isolate the first column of $Enc[Z]$ and exclude the other columns, we can create a constant matrix $F$ with ones in the first column and zeros elsewhere. Multiplying $Enc[Z]$ by $F$ will yield the desired ciphertext.
Han et al. (2019) mentioned several basic but important operations used by Kim et al. (2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix. With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
2.3 Logistic Regression Model
Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not. LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\boldsymbol{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\boldsymbol{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate. Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
$\Pr(y=+1|\mathbf{x},\boldsymbol{\beta})$ $=\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\boldsymbol{\beta}^{\top}\mathbf{x}}},$
$\Pr(y=-1|\mathbf{x},\boldsymbol{\beta})$ $=1-\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\boldsymbol{\beta}^{\top}\mathbf{x}}}.$
LR sets a threshold (usually $0.5$) and compares its output with it to decide the resulting class label.
The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\boldsymbol{\beta}$ to maximize $L(\boldsymbol{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\boldsymbol{\beta})$ or its log-likelihood function $l(\boldsymbol{\beta})$ for convenience in the calculation:
$l(\boldsymbol{\beta})=\ln L(\boldsymbol{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}}),$
where $n$ is the number of examples in the training dataset. LR does not have a closed form of maximizing $l(\boldsymbol{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix. The gradient and Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ are given by, respectively:
$\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
$\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
$=X^{\top}SX,$
where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero. Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\boldsymbol{\beta}$ in $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})=0$ for LR.
3 Technical Details
It is quite time-consuming to compute the Hessian matrix and its inverse in Newton’s method for each iteration. One way to limit this downside is to replace the varying Hessian with a fixed matrix $\bar{H}$. This novel technique is called the fixed Hessian Newton’s method. Böhning and Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is non-negative definite. With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\bar{H}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}).$
Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ in LR.
3.1 Simplified Fixed Hessian
Bonte and Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain. They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$. They also suggest a specific order of calculation to optimize the computation of $B$ more efficiently. Their new approximation $B$ of the fixed Hessian is:
$B=\left[\begin{array}[]{cccc}\sum_{i=0}^{d}\bar{h}_{0i}&0&\ldots&0\\ 0&\sum_{i=0}^{d}\bar{h}_{1i}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\sum_{i=0}^{d}\bar{h}_{di}\\ \end{array}\right],$
where $\bar{h}_{ki}$ is the element of $\bar{H}$. This diagonal matrix $B$ is in a very simple form and can be obtained from $\bar{H}$ without much difficulty. The inverse of $B$ can even be approximated in the encrypted form by computing the inverse of each diagonal element of $B$ using an iterative Newton’s method with an appropriate initial value. Their simplified fixed Hessian method can be formulated as follows:
$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-B^{-1}\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]$
$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$.
Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $-\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\boldsymbol{\beta}_{t}+\eta\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
which is the same as the formula of the naive gradient $ascent$ method. Such a coincidence not only helps generate the idea behind this work but also leads us to believe that there is a connection between the Hessian matrix and the learning rate of the gradient (descent) method.
We regard $B^{-1}\cdot\nabla_{i}$ as a novel enhanced gradient variant and allocate a distinct learning rate to it. As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
The SFH method proposed by Bonte and Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive. For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method. However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets. We extend their work by removing these limitations so as to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term quadratic gradient.
3.2 Quadratic Gradient Definition
Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\boldsymbol{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
$\boldsymbol{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
$\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$. We construct a new diagnoal Hessian matrix $\tilde{B}$ with each diagnoal element $\tilde{B}_{kk}$ being $-\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$,
$\tilde{B}=\left[\begin{array}[]{ccc}-\epsilon-\sum_{i=0}^{d}|\bar{h}_{0i}|&0&\ldots\\ 0&-\epsilon-\sum_{i=0}^{d}|\bar{h}_{1i}|&\ldots\\ \vdots&\vdots&\ddots\\ 0&0&\ldots\\ \end{array}\right],$
where $\epsilon$ is a small positive constant to avoid division by zero (usually set to $81e-$ 8).
As long as $\tilde{B}$ satisfies the convergence condition of the above fixed Hessian method, $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound. Since we already assume that $\bar{H}\leq H$, it will suffice to show that $\tilde{B}\leq\bar{H}$. We prove $\tilde{B}\leq\bar{H}$ in a similar way that (Bonte & Vercauteren, 2018) did.
Lemma 3.1.
Let $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix, and let $B$ be the diagonal matrix whose diagonal entries $B_{kk}=-\epsilon-\sum_{i}|A_{ki}|$ for $k=1,\ldots,n$, then $B\leq A$.
Proof.
By definition of the Loewner ordering, we have to prove the difference matrix $C=A-B$ is non-negative definite, which means that all the eigenvalues of $C$ need to be non-negative. By construction of $C$ we have that $C_{ij}=A_{ij}+\epsilon+\sum_{k=1}^{n}|A_{ik}|$ for $i=j$ and $C_{ij}=A_{ij}$ for $i\neq j$. By means of Gerschgorin’s circle theorem, we can bound every eigenvalue $\lambda$ of $C$ in the sense that $|\lambda-C_{ii}|\leq\sum_{i\neq j}|C_{ij}|$ for some index $i\in\{1,2,\ldots,n\}$. We conclude that $\lambda\geq A_{ii}+\epsilon+|A_{ii}|\geq\epsilon>0$ for all eigenvalues $\lambda$ and thus that $B\leq A$. ∎
Definition 3.2 (Quadratic Gradient).
Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\boldsymbol{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps. Note that $G$ is still a column vector of the same size as the gradient $\boldsymbol{g}$. To maximize the function $F(\mathbf{x})$, we can use the iterative formulas: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+\eta\cdot G$, just like the naive gradient.
To minimize the function $F(x)$ is the same as to just maximize the function $-F(x)$, in which case we need to construct the $\tilde{B}$ by any good lower bound $\bar{H}$ of the Hessian $-H$ of $-F(x)$ or any good upper bound $\bar{H}$ of the Hessian $H$ of $F(x)$. We point out here that $\bar{H}$ could be the Hessian matrix $H$ itself.
Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
1:  Require: training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
2:  Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
5:  for $i:=0$ to $d$ do
6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
7:     $\bar{B}[i][i]\leftarrow\epsilon$
8:     for $j:=0$ to $d$ do
9:        $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
10:     end for
11:  end for
12:  Set $alpha_{0}\leftarrow 0.01$
13:  Set $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
14:  for $count:=1$ to $\kappa$ do
15:     # $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
16:     Set $Z\leftarrow\boldsymbol{0}$
17:     for $i:=1$ to $n$ do
18:        for $j:=0$ to $d$ do
19:           $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
20:        end for
21:     end for
22:     # $\boldsymbol{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
23:     Set $\boldsymbol{\sigma}\leftarrow\boldsymbol{0}$
24:     for $i:=1$ to $n$ do
25:        $\boldsymbol{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
26:     end for
27:     Set $\boldsymbol{g}\leftarrow\boldsymbol{0}$
28:     for $j:=0$ to $d$ do
29:        for $i:=1$ to $n$ do
30:           $\boldsymbol{g}[j]\leftarrow\boldsymbol{g}[j]+(1-\boldsymbol{\sigma}[i])\times Y[i]\times X[i][j]$
31:        end for
32:     end for
33:     Set $G\leftarrow\boldsymbol{0}$
34:     for $j:=0$ to $d$ do
35:        $G[j]\leftarrow\bar{B}[j][j]\times\boldsymbol{g}[j]$
36:     end for
37:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
38:     Set $\gamma\leftarrow lr/(n\times count)$
39:     for $j:=0$ to $d$ do
40:        $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
41:        $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
42:        $W[j]\leftarrow w_{temp}$
43:     end for
44:     $alpha_{0}\leftarrow alpha_{1}$
45:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
46:  end for
47:  Return $V$
3.3 Quadratic Gradient Algorithms
This gradient variant, Quadratic Gradient, can be used to enhance various first-order gradient algorithms:
  1. 1.
    NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient $ascent$ method for NAG are as follows:
    $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
    $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a). The enhanced NAG is to replace (3) with $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually setted to $1+\eta_{t}$. Our enhanced NAG is described in Algorithm 1 .
  2. 2.
    Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\boldsymbol{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
  3. 3.
    Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively. Similar to how Adagrad transitions to its enhanced versions, the update formula for enhanced Adam, also employs a new learning rate and replaces the gradient with its quadratic form.
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 1: The training results of NAG and Enhanced NAG in the clear
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 3: The training results of Adam and Enhanced Adam in the clear
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function, specifically maximum likelihood estimation (MLE), as the only indicator. We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al. (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (Edinburgh), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work (Kim et al., 2018a), we use the learning rate $1+\frac{10}{1+t}$, in line with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the binary logistic regression model.
Figures 21, and 3 show that, except for the Enhanced Adagrad and Adam methods on the iDASH genomic dataset, our enhanced methods converge faster than their original counterparts in all other cases. All figures demonstrate that the enhanced algorithms converge empirically. In all the Python experiments, the time required to calculate $\bar{B}$ in the quadratic gradient $G$ before running the iterations, as well as the time to execute each iteration for various algorithms, is minimal and can be considered negligible (a few seconds).
An important observation is that the enhanced algorithms have a better performance when using learning rates between $1$ and $2$ than using learning rate otherelse. When using a learning rate greater than $3$, the algorithm is almost certain not to converge. When employing quadratic gradient algorithms, we recommend adopting the learning rate setting $1+A\cdot\gamma^{t}$ as the optimal choice, where $t$ represents the iteration number, $A$ is a positive float number typically set to no less than $1$, and $\gamma$ is usually a positive number less than 1, controlling the rate of decay. We strongly recommend that $A$ is set to $1.0$ and $\gamma$ is tuned to search for optimal results, or that one can set $A$ to a small number less than $1.0$ and $\gamma$ to a larger number greater than $0.9$.
Analysis In Figure 2(a), the enhanced Adagrad algorithm failed to outperform the original Adagrad algorithm. The possible reason for that might be related to the limitations of the raw Adagrad method. Without a doubt, Adagrad is a novel algorithm initiated to accelerate each element of the gradient with different learning rates. However, Adagrad tends to converge to a suboptimal solution due to its aggressive, monotonically decreasing learning rates. This would lead to its main limitation that in the later training phase every learning rate for different components of the gradient is too close to zero due to keeping adding positive additional terms to the denominator, stopping the algorithm from learning anything.
On the other hand, the original Adagrad method has another little-noticed limitation: the learning rate in the first few iterations tends to be large. While this limitation does not affect the performance of the original Adagrad method to some extent, the enhanced Adagrad method exacerbates this phenomenon significantly, leading to the Learning-Rate Explosion . Therefore, the enhanced Adagrad cannot be applied to general optimization problems such as Rosenbrock’s function. The exploding learning rate would be too large for the algorithm to survive the first several iterations, finally leading the optimization function to some point where its output cannot be represented by the computer system. This might explain why the performance of this algorithm in all cases, not just on the iDASH genome dataset, seems to be meaningless and unstable in the first few iterations.
Several improved algorithms upon the Adagrad method, such as RMSProp, have been proposed in order to address these issues existed, via using an exponential moving average of historical gradients rather than just the sum of all squared gradients from the beginning of training. We might be able to overcome the problems existing in the enhanced Adagrad method by adopting the enhanced Adagrad-like variants, like the enhanced Adadelta method and the enhanced RMSProp method.
The reasons why the enhanced algorithms have a better performance might be that quadratic gradient can incorporate curves into first-order gradient algorithms.
(a) The raw first-order gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset
Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and raw quadratic gradient algorithm on the lbw dataset using various learning rates. Figure 4 shows the detailed results of this experiment. It is precisely these results that inspired the present authors to come up with the idea of quadratic gradient. Similar to the gradient, the quadratic gradient carrying the curve also exhibits the characteristic of smooth progression. That is, when the learning rate is gradually increased, both the gradient and the quadratic gradient will show corresponding gradual changes in performance rather than abrupt shifts.
4 Homomorphic Training
Adagrad method is not a practical solution for homomorphic LR due to its frequent inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
X $=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$
$\text{ct}_{Z}$ $=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
$\text{ct}_{\beta}^{(0)}$ $=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$
$\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$. Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.
Limitations
In a privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
Table 1: Implementation Results for iDASH datasets with 10-fold CV
Dataset
Sample
Num
Feature
Num
Method deg $g$
Iter
Num
Storage
Learn
Time
Accuracy
AUC
iDASH 1579 18 Ours 5 4 0.08 4.43 61.46 0.696
(Kim et al., 2018a) 5 7 0.04 6.07 62.87 0.689
Table 2: Implementation Results for other datasets with 5-fold CV
Dataset
Sample
Num
Feature
Num
Method deg $g$
Iter
Num
Storage
Learn
Time
Accuracy
AUC
Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
(Kim et al., 2018a) 5 7 0.02 3.6 91.04 0.958
lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
(Kim et al., 2018a) 5 7 0.02 3.3 69.19 0.689
nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
(Kim et al., 2018a) 5 7 0.16 7.3 79.22 0.717
pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
(Kim et al., 2018a) 5 7 0.02 3.5 68.27 0.740
uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
(Kim et al., 2018a) 5 7 0.02 3.5 74.44 0.603
5 Experiments
Implementation
We implement the enhanced NAG based on HE with the library HEAAN. The C++ source code is publicly available at https://anonymous.4open.science/r/IDASH2017-245B . All the experiments on the ciphertexts were conducted on a public cloud with 32 vCPUs and 64 GB RAM.
For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
The parameters of HEAAN we set are same to (Kim et al., 2018a): $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to (Kim et al., 2018a) for details on these parameters. Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method. Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result.
6 Conclusion
In this paper, we proposed a faster gradient variant called quadratic gradient, and implemented the quadratic-gradient version of NAG in the encrypted domain to train the logistic regression model.
The quadratic gradient presented in this work can be constructed from the Hessian matrix directly, and thus somehow combines the first-order Newton’s method and the second-order gradient (descent) method together. There is a promising chance that quadratic gradient could accelerate other gradient methods, which is an open future work.
Also, quadratic gradient might substitute and supersede the line-search method, for example when using enhanced Adagrad-like methods, and could use gradient descent methods to accelerate Newton’s method, resulting in super-quadratic algorithms.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.