### v13

# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

Abstract
Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years. In this paper, we introduce an efficient gradient variant, called $quadratic$ $gradient$, which can be used for privacy-preserving logistic regression training. We enhance Nesterov’s Accelerated Gradient (NAG), Adaptive Gradient Algorithm (Adagrad) and Adam algorithms by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets. Experimental results demonstrate that the enhanced algorithms achieve significantly improved convergence speed compared to traditional first-order gradient methods. Moreover, we applied the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results within just 4 iterations. There is a great chance that the quadratic gradient approach could integrate first-order gradient descent/ascent algorithms with the second-order Newton-Raphson methods, and that quadratic gradient could be applied to a wide range of numerical optimization problems.
Index Terms:
Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
I Introduction
Given a person’s healthcare data related to a certain disease, we can train a logistic regression (LR) model capable of telling whether or not this person is likely to develop this disease. However, such personal health information is highly private to individuals. The privacy concern, therefore, becomes a major obstacle for individuals to share their biomedical data. The most secure solution is to encrypt the data into ciphertexts first by Homomorphic Encryption (HE) and then securely outsource the ciphertexts to the cloud, without allowing the cloud to access the data directly. iDASH is an annual competition that aims to call for implementing interesting cryptographic schemes in a biological context. Since 2014, iDASH has included the theme of genomics and biomedical privacy. The third track of the 2017 iDASH competition and the second track of the 2018 iDASH competition were both to develop HE-based solutions for building an LR model over encrypted data.
Several studies on logistic regression models are based on homomorphic encryption. Kim et al. [1] discussed the problem of performing LR training in an encrypted environment. They employed full batch gradient descent in the training process and utilized the least-squares method to approximate the sigmoid function. In the 2017 iDASH competition, Bonte and Vercauteren [2], Kim et al. [3], Chen et al. [4], and Crawford et al. [5] all investigated the same problem that Kim et al. [1] studied. In the iDASH competition of 2018, Kim et al. [6] and Blatt et al. [7] further worked on it for an efficient packing and semi-parallel algorithm. There are other related works [8, 9] focusing on various aspects but the papers most relevant to this work are [2] and [3]. Bonte and Vercauteren [2] developed a practical algorithm called the simplified fixed Hessian (SFH) method. Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al. [3] for efficient homomorphic computation.
Our specific contributions in this paper are as follows:
  1. 1.
    We introduce a novel gradient variant, termed quadratic gradient, that bridges first-order gradient-based algorithms and the second-order Newton–Raphson (i.e., Newton’s) method, thereby enabling a unified framework that leverages the strengths of both approaches.
  2. 2.
    We develop three enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms exhibit state-of-the-art performance in terms of convergence speed.
  3. 3.
    We implement privacy-preserving logistic regression training using the enhanced Nesterov’s Accelerated Gradient (NAG) method, which, to the best of our knowledge, provides an effective balance between computation and storage without significant compromise.
  4. 4.
    We propose a systematic framework to determine the existence of constant Hessian approximations in fixed-Hessian methods [10, 2], and to assess the feasibility of employing fixed Hessian matrices for efficient computation.
II Preliminaries
We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows. For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
II-A Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation [11]. FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext [12] also contributes to the slowdown. Cheon et al. [13] proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in [3, 1, 14].
II-B Database Encoding Method
Kim et al. [3] proposed an efficient and promising database encoding method leveraging the SIMD technique, which enables effective utilization of both computational and storage resources. Given a training dataset $Z$ consisting of $n$ samples with $(1+d)$ covariates each, the authors packed $Z$ into a single ciphertext in a row-wise manner, thereby maximizing the use of available ciphertext slots.
When employing this encoding scheme, we can manipulate the data matrix $Z$ through HE operations on the ciphertext $Enc[Z]$, utilizing only three HE operations - rotation, addition, and multiplication. For instance, if we wish to isolate the first column of $Enc[Z]$ and exclude the other columns, we can create a constant matrix $F$ with ones in the first column and zeros elsewhere. Multiplying $Enc[Z]$ by $F$ will yield the desired ciphertext.
Han et al. [14] mentioned several basic but important operations used by Kim et al. [3] in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix. With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
II-C Logistic Regression Model
Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not. LR can be generalized from linear regression [15] by mapping the whole real line $(\boldsymbol{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\boldsymbol{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate. Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
$\Pr(y=+1|\mathbf{x},\boldsymbol{\beta})$ $=\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\boldsymbol{\beta}^{\top}\mathbf{x}}},$
$\Pr(y=-1|\mathbf{x},\boldsymbol{\beta})$ $=1-\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\boldsymbol{\beta}^{\top}\mathbf{x}}}.$
LR sets a threshold (usually $0.5$) and compares its output with it to decide the resulting class label.
The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\boldsymbol{\beta}$ to maximize $L(\boldsymbol{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\boldsymbol{\beta})$ or its log-likelihood function $l(\boldsymbol{\beta})$ for convenience in the calculation:
$l(\boldsymbol{\beta})=\ln L(\boldsymbol{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}}),$
where $n$ is the number of examples in the training dataset. LR does not have a closed form of maximizing $l(\boldsymbol{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix. The gradient and Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ are given by, respectively:
$\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
$\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
$=X^{\top}SX,$
where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum [16], where its gradient is zero. Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\boldsymbol{\beta}$ in $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})=0$ for LR.
III Technical Details
Newton’s method requires computing the Hessian matrix and its inverse at every iteration, which is computationally intensive. To address this, the fixed Hessian Newton’s method (also known as the lower bound principle) replaces the varying Hessian with a constant matrix $\bar{H}$, offering improved efficiency. Böhning and Lindsay [10] have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is non-negative definite. With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\bar{H}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}).$
Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ in LR.
III-A Simplified Fixed Hessian
Bonte and Vercauteren [2] further simplified the fixed Hessian $\bar{H}$ to facilitate its inversion in the encrypted domain. They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$. They also suggest a specific order of calculation to optimize the computation of $B$ more efficiently. Their new approximation $B$ of the fixed Hessian is:
$B=\left[\begin{array}[]{cccc}\sum_{i=0}^{d}\bar{h}_{0i}&0&\ldots&0\\ 0&\sum_{i=0}^{d}\bar{h}_{1i}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\sum_{i=0}^{d}\bar{h}_{di}\\ \end{array}\right],$
where $\bar{h}_{ki}$ is the element of $\bar{H}$. This diagonal matrix $B$ is in a very simple form and can be obtained from $\bar{H}$ without much difficulty. The inverse of $B$ can even be approximated in the encrypted form by computing the inverse of each diagonal element of $B$ using an iterative Newton’s method with an appropriate initial value. Their simplified fixed Hessian method can be formulated as follows:
$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-B^{-1}\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]$
$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$.
Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $-\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\boldsymbol{\beta}_{t}+\eta\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
which is the same as the formula of the naive gradient $ascent$ method. Such a coincidence not only helps generate the idea behind this work but also leads us to believe that there is a connection between the Hessian matrix and the learning rate of the gradient (descent) method.
We introduce $B^{-1}\cdot\nabla_{\boldsymbol{\beta}}\ell(\boldsymbol{\beta})$ as a novel enhanced gradient variant, and allocate a dedicated learning rate to better capture its dynamics. As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
The SFH method proposed by Bonte and Vercauteren [2] has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive. For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method. However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets. We extend their work by removing these limitations so as to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term quadratic gradient.
III-B Quadratic Gradient: Definition
Suppose a differentiable scalar-valued function $F(\mathbf{x})$ has gradient $\boldsymbol{g}$ and Hessian matrix $H$. For the $maximization$ problem, let $\bar{H}$ be any matrix satisfying $\bar{H}\preceq H$ in the Loewner partial ordering. We represent these quantities as:
$\boldsymbol{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
$\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$. We construct a new diagonal matrix $\tilde{B}$, referred to as the diagonal Hessian approximation, where each diagonal entry $\tilde{B}_{kk}$ is defined as $-\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$. Formally, $\tilde{B}$ is given by:
$\tilde{B}=\left[\begin{array}[]{ccc}-\epsilon-\sum_{i=0}^{d}|\bar{h}_{0i}|&0&\ldots\\ 0&-\epsilon-\sum_{i=0}^{d}|\bar{h}_{1i}|&\ldots\\ \vdots&\vdots&\ddots\\ 0&0&\ldots\\ \end{array}\right],$
where $\epsilon>0$ is a small constant (typically set to $1\text{e}{-8}$) added for numerical stability, such as preventing division by zero.
As long as $\tilde{B}$ satisfies the convergence condition of the above fixed Hessian method, $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound. Since we already assume that $\bar{H}\leq H$, it will suffice to show that $\tilde{B}\leq\bar{H}$. We prove $\tilde{B}\leq\bar{H}$ in a similar way that [2] did.
Lemma III.1.
Let $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix, and let $B$ be the diagonal matrix whose diagonal entries $B_{kk}=-\epsilon-\sum_{i}|A_{ki}|$ for $k=1,\ldots,n$, then $B\leq A$.
Proof.
By definition of the Loewner ordering, we have to prove the difference matrix $C=A-B$ is non-negative definite, which means that all the eigenvalues of $C$ need to be non-negative. By construction of $C$ we have that $C_{ij}=A_{ij}+\epsilon+\sum_{k=1}^{n}|A_{ik}|$ for $i=j$ and $C_{ij}=A_{ij}$ for $i\neq j$. By means of Gerschgorin’s circle theorem, we can bound every eigenvalue $\lambda$ of $C$ in the sense that $|\lambda-C_{ii}|\leq\sum_{i\neq j}|C_{ij}|$ for some index $i\in\{1,2,\ldots,n\}$. We conclude that $\lambda\geq A_{ii}+\epsilon+|A_{ii}|\geq\epsilon>0$ for all eigenvalues $\lambda$ and thus that $B\leq A$. ∎
Definition III.2 (Quadratic Gradient).
Given the previously defined $\tilde{B}$, we define the quadratic gradient as $G=\bar{B}\cdot\boldsymbol{g}$ with a new learning rate $N_{t}$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$. The learning rate $N_{t}$ should always be no less than 1 and is designed to decrease to 1 over a limited number of iterations. It is important to note that $G$ remains a column vector of the same dimension as the gradient $\boldsymbol{g}$. To maximize the function $F(\mathbf{x})$, we can employ the iterative formula: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+N_{t}\cdot G$, similar to the naive gradient approach.
Minimizing the function $F(x)$ is equivalent to maximizing the function $-F(x)$. In this context, we need to construct $\tilde{B}$ using either a good $lower$ bound $\bar{H}$ for the Hessian $-H$ of $-F(x)$ or a good $upper$ bound $\bar{H}$ for the Hessian $H$ of $F(x)$. It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
Böhning and Lindsay [10] did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method. This may be due to the absence of such matrices for most objective functions. In this work, we present a systematic approach for identifying if there were fixed Hessian matrices existing. Our proposed method serves as one of many possible approaches to detect the existence of a fixed Hessian matrix. Specifically, if our approach identifies a fixed Hessian, this provides sufficient evidence for the existence of such a matrix. However, failure to detect a fixed Hessian via our solution does not imply its non-existence. In other words, detection by our technique constitutes a sufficient but not necessary condition for the existence of a fixed Hessian matrix.
We note that our quadratic gradient approach does not require a constant (fixed) Hessian replacement and instead allows for the use of the original Hessian itself to construct our varying diagonal matrix $\tilde{B}$ for the quadratic gradient. Under this less restrictive condition, we present a methodical approach to evaluate whether a fixed Hessian matrix can be employed for efficient computation.
Specifically, our method proceeds in the following steps:
  • Step 1: Construct $\tilde{B}$, an approximation of the Hessian, directly from the Hessian of the given objective function.
  • Step 2: For each diagonal element $\tilde{B}_{kk}$, examine whether there exists a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, or equivalently, a constant minimum of $|\tilde{B}_{kk}|$. If each diagonal element $|\tilde{B}_{kk}|$ achieves its minimum $|\tilde{M}_{kk}|$, then a constant diagonal matrix with the corresponding diagonal elements $1/|\tilde{M}_{kk}|$ would satisfy the convergence condition for the SFH method.
  • Step 3: Evaluate the performance of the resulting fixed Hessian approximation. While this replacement satisfies theoretical convergence conditions, it may not always yield a tight bound to the true Hessian, potentially degrading the convergence rate in practice.
This search method can be applied to obtain the same fixed Hessian matrix $-\frac{1}{4}X^{\top}X$ as originally proposed by the authors in [10], in the context of binary logistic regression.
Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm
1:  Require: training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
2:  Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
5:  for $i:=0$ to $d$ do
6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
7:     $\bar{B}[i][i]\leftarrow\epsilon$
8:     for $j:=0$ to $d$ do
9:        $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
10:     end for
11:  end for
12:  Set $alpha_{0}\leftarrow 0.01$
13:  Set $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
14:  for $count:=1$ to $\kappa$ do
15:     # $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
16:     Set $Z\leftarrow\boldsymbol{0}$
17:     for $i:=1$ to $n$ do
18:        for $j:=0$ to $d$ do
19:           $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
20:        end for
21:     end for
22:     # $\boldsymbol{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
23:     Set $\boldsymbol{\sigma}\leftarrow\boldsymbol{0}$
24:     for $i:=1$ to $n$ do
25:        $\boldsymbol{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
26:     end for
27:     Set $\boldsymbol{g}\leftarrow\boldsymbol{0}$
28:     for $j:=0$ to $d$ do
29:        for $i:=1$ to $n$ do
30:           $\boldsymbol{g}[j]\leftarrow\boldsymbol{g}[j]+(1-\boldsymbol{\sigma}[i])\times Y[i]\times X[i][j]$
31:        end for
32:     end for
33:     Set $G\leftarrow\boldsymbol{0}$
34:     for $j:=0$ to $d$ do
35:        $G[j]\leftarrow\bar{B}[j][j]\times\boldsymbol{g}[j]$
36:     end for
37:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
38:     Set $\gamma\leftarrow lr/(n\times count)$
39:     for $j:=0$ to $d$ do
40:        $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
41:        $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
42:        $W[j]\leftarrow w_{temp}$
43:     end for
44:     $alpha_{0}\leftarrow alpha_{1}$
45:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
46:  end for
47:  Return $V$
III-C Quadratic Gradient: Algorithms
Quadratic gradient can be used to enhance various first-order gradient algorithms:
  1. 1.
    NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient $ascent$ method for NAG are as follows:
    $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
    $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position [3]. The enhanced NAG algorithm modifies the original update rule (3) by replacing it with
    $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot B^{-1}\cdot\nabla J(\boldsymbol{\beta}_{t})=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$
    where $N_{t}$ denotes a learning rate adapted specifically for the enhanced variant. As detailed in Algorithm 1, we typically set $N_{t}=1+\eta_{t}$, where $\eta_{t}$ follows a decay schedule or fixed setting depending on the application.
  2. 2.
    Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\boldsymbol{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
  3. 3.
    Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively. Similar to how Adagrad transitions to its enhanced version, the update formula for enhanced Adam, also employs a new learning rate and replaces the gradient with its quadratic form.
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 1: The training results of NAG vs. Enhanced NAG in the clear domain.
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 2: The training results of Adagrad vs. Enhanced Adagrad in the clear domain.
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 3: The training results of Adam vs. Enhanced Adam in the clear domain.
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\boldsymbol{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator. We conduct a comprehensive evaluation of six optimization algorithms—namely, NAG, Adagrad, Adam, and their respective quadratic-gradient enhanced variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam)—on the benchmark datasets utilized by [3], including the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), NHANES III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from [14]: a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work by [3], the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. The Enhanced Adagrad algorithm employs a modified learning rate of $N_{t}=0.06+0.01$, in contrast to the default setting $N_{t}=0.01$ used by the original Adagrad algorithm. Similarly, the Enhanced Adam algorithm adopts an adjusted learning rate configuration with $\alpha=0.01+0.001$ (the new learning rate), $\beta_{1}=0.9$, and $\beta_{2}=0.999$, whereas the standard Adam algorithm follows its default settings of $\alpha=0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. For further details regarding these parameters, refer to [17]. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\boldsymbol{\beta})$.
From the empirical results presented in Figures 21, and 3, we conclude that all the enhanced algorithms converge and consistently outperforms the original NAG method in terms of convergence rate.
An important observation is that the enhanced algorithms demonstrate better performance with learning rates between $1$ and $2$ compared to other values. When the learning rate exceeds $3$, the algorithm is nearly guaranteed not to converge. For quadratic-gradient NAG algorithms, we recommend employing an exponentially decaying learning rate, such as one of the form $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration index, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive scalar less than $1$ that governs the rate of decay.
(a) First-order gradient ascent: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
(b) Quadratic gradient ascent: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
Figure 4: Training results of standard first-order and quadratic gradient ascent algorithms on the lbw dataset in the plaintext domain.
Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and raw quadratic gradient algorithm on the lbw dataset using various learning rates. Figure 4 shows the detailed results of this experiment. It is precisely these results that inspired the present authors to come up with the idea of quadratic gradient. Similar to the gradient, the quadratic gradient also exhibits the characteristic of smooth progression. That is, when the learning rate is gradually increased, both the gradient and the quadratic gradient will show corresponding gradual changes in performance rather than abrupt shifts.
IV Secure Training
Both Adagrad and Adam algorithms are not practical solutions for homomorphic logistic regression due to their frequent reliance on inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. [3], utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. [3] used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The initial weight vector $\beta^{(0)}$, initialized as a zero vector, and the diagonal elements of the matrix $\bar{B}$ are each replicated $n$ times to construct two corresponding matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
X $=\left[\begin{array}[]{cccc}1&x_{11}&\ldots&x_{1d}\\ 1&x_{21}&\ldots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\ldots&x_{nd}\\ \end{array}\right],\text{Y}=\left[\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\\ \end{array}\right],$
$\text{ct}_{Z}$ $=Enc\left[\begin{array}[]{cccc}y_{1}&y_{1}x_{11}&\ldots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\ldots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\ldots&y_{n}x_{nd}\\ \end{array}\right],$
$\text{ct}_{\beta}^{(0)}$ $=Enc\left[\begin{array}[]{cccc}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\ldots&\beta_{d}^{(0)}\\ \end{array}\right],$
$\text{ct}_{\bar{B}}$ $=Enc\left[\begin{array}[]{cccc}\bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{[0][0]}&\bar{B}_{[1][1]}&\ldots&\bar{B}_{[d][d]}\\ \end{array}\right],$
where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
The public cloud receives the ciphertexts $\text{ct}Z$, $\text{ct}{\beta}^{(0)}$, and $\text{ct}{\bar{B}}$, and executes the enhanced NAG algorithm to iteratively refine the encrypted weight vector by updating $\text{ct}{\beta}^{(0)}$ toward a more optimal solution. Refer to [3] for a detailed description about how to calculate the gradient by HE programming.
Limitations
In the privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.
TABLE I: Implementation Results for iDASH datasets with 10-fold CV
Dataset
Sample
Num
Feature
Num
Method deg $g$
Iter
Num
Storage
Learn
Time
Accuracy
AUC
iDASH 1579 18 Ours 5 4 0.08 4.43 61.46 0.696
[3] 5 7 0.04 6.07 62.87 0.689
TABLE II: Implementation Results for other datasets with 5-fold CV
Dataset
Sample
Num
Feature
Num
Method deg $g$
Iter
Num
Storage
Learn
Time
Accuracy
AUC
Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
[3] 5 7 0.02 3.6 91.04 0.958
lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
[3] 5 7 0.02 3.3 69.19 0.689
nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
[3] 5 7 0.16 7.3 79.22 0.717
pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
[3] 5 7 0.02 3.5 68.27 0.740
uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
[3] 5 7 0.02 3.5 74.44 0.603
V Experiments
We implement the enhanced NAG algorithm over homomorphically encrypted data using the HEAAN library. The C++ source code is publicly available at https://anonymous.4open.science/r/IDASH2017-245B. All experiments on encrypted data were conducted on a public cloud instance equipped with 32 vCPUs and 64 GB of RAM.
For a fair comparison with the baseline in [3], we adopt the same 10-fold cross-validation (CV) technique on the iDASH dataset, which contains 1,579 samples with 18 features, and apply the same 5-fold CV setup to the remaining five datasets. Following [3], we report average accuracy and Area Under the Curve (AUC) as the primary evaluation metrics. The results of these experiments are summarized in Tables I and II, which also include average evaluation runtime and storage consumption (i.e., encrypted dataset for the baseline, and encrypted dataset plus $\bar{B}$ for our method). To ensure consistency in ciphertext storage, we adopt the same packing strategy proposed by Kim et al. [3], resulting in comparable ciphertext sizes, with the only addition being the ciphertexts required to encode $\bar{B}$. We use a learning rate of $1+0.9^{t}$ throughout all training procedures.
The parameters for HEAAN are configured identically to those in [3], with $\log N=16$, $\log Q=1200$, $\log p=30$, and the number of slots set to $32{,}768$, yielding a security level of $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to [3] for details on these parameters. Since our enhanced NAG method requires one additional ciphertext multiplication compared to the baseline approach—resulting in greater modulus consumption—our scheme is limited to performing only four iterations of the enhanced NAG algorithm. Nevertheless, even with just four iterations, the proposed method achieves results that remain comparable to those of the baseline. Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for encryption and evaluation. It is worth noting that we relied on their open-source code to implement our homomorphic logistic regression training. Therefore, theoretically, the running time for one iteration of our algorithm should be very similar to theirs.
VI Conclusion
In this paper, we introduce a novel gradient variant, referred to as the quadratic gradient, and implement a quadratic-gradient-enhanced version of Nesterov’s Accelerated Gradient (NAG) to train a logistic regression model in the encrypted domain. The proposed quadratic gradient is preferably constructed directly from the Hessian matrix, effectively bridging first-order gradient-based methods with second-order Newton-type techniques. This approach exhibits strong potential for accelerating a broad range of optimization algorithms. In particular, it offers a viable alternative to traditional line-search procedures, can be incorporated into adaptive methods such as enhanced AdaGrad, and may serve to accelerate Newton’s method via embedded gradient descent steps.
As future work, we aim to establish a more rigorous theoretical analysis of the convergence properties of the enhanced algorithm. A more practical and application-driven direction involves systematically evaluating the performance of the enhanced method under different learning rate configurations.